Rationalize the denominator:

2 / √[3] + √[2]

Answer by malakumar_kos@yahoo.com(315)


2 / √[3] + √[2]
rationalising factor of sq rt3+ sq rt 2 = sq rt 3- sq rt 2
multiply and divide by sq rt 3 - sq rt 2
2(sq rt 3- sq rt 2)/( sq rt 3+ sq rt 2.)(sq rt 3- sq rt 2)
= 2( sq rt 3- sq rt 2)/(sq rt 3)^2- (sq rt 2)^2
= 2(sq rt 3- sq rt 2)/3-2
= 2(sq rt 3- sq rt 2)/1
= 2(sq rt 3- sq rt 2)